what is 3x^2 + 3y^2 + 24y + 18x - 27=0 written in standard form
Answer by stanbon(75887)

what is 3x^2 + 3y^2 + 24y + 18x - 27=0 written in standard form
----
Gather the x-terms; Gather the y-terms:
3x^2 +18x + 3y^2 + 24y = 27
-----
Divide thru by 3 to get:
x^2 + 6x + y^2 + 8y = 9
----
complete the square on the x-terms and on the y-terms:
x^2 + 6x + 9 + y^2 + 8y + 16 = 9 + 9 + 16
----
Factor:
(x+3)^2 + (y+4)^2 = 34
That is the standard form:
==============================
cheers,
Stan H.
===============================